# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v1

@@ A.2 Experimental Setting Information
Non-Linear Bandit: We now consider the setting when $f(\mathbf{x},\bm{\theta}_{*})=1/(1+0.5\cdot\exp(2\cdot\exp(-\mathbf{x}^{\top}\bm{\theta}_{*})))$. Again, here $\mathbf{x}\in\mathbb{R}^{d}$ is the action feature, and $\bm{\theta}_{*}\in\mathbb{R}^{d}$ is the hidden parameter. Note that this is different than the generalized linear bandit setting [Filippi et al., 2010, Li et al., 2017]. Again for every experiment, we first generate tasks from $\mathcal{T}_{\text{pre}}$. Then we sample a fixed set of actions from $\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}/d\right)$ in $\mathbb{R}^{d}$ and this constitutes the features. Then for each task $m\in[M]$ we sample $\bm{\theta}_{m,*}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}/d\right)$ to produce the means $\mu(m,a)=1/(1+0.5\cdot\exp(2\cdot\exp(-\mathbf{x}(m,a)^{\top}\bm{\theta}_{m,*})))$ for $a\in\mathcal{A}$ and $m\in[M]$. Again note that in this setting $\mathbf{x}(m,a)$ for each $a\in\mathcal{A}$ is fixed for all tasks $m$.
@@
A.3 Empirical Study: Bilinear Bandits
@@
In this section, we discuss the performance of PreDeToR against the other baselines in the bilinear setting. Again note that the number of tasks $M_{\text{pre}}\gg A\geq n$. Through this experiment, we want to evaluate the performance of PreDeToR to exploit the underlying latent structure and reward correlation when the horizon is small, the number of tasks is large, and understand its performance in the bilinear bandit setting [Jun et al., 2019, Lu et al., 2021, Kang et al., 2022, Mukherjee et al., 2023]. Note that this setting also goes beyond the linear feedback model [Abbasi-Yadkori et al., 2011, Lattimore and Szepesvári, 2020] and is related to matrix bandits [Yang and Wang, 2020].
@@
This special case is studied in Chaudhuri et al. [2017].
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, and Thomp. However, we now implement the LowOFUL [Jun et al., 2019] which is optimal in the bilinear bandit setting.
LowOFUL: The LowOFUL algorithm first estimates the unknown parameter $\mathbf{\Theta}_{m,*}$ for each task $m$ using E-optimal design [Pukelsheim, 2006, Fedorov, 2013, Jun et al., 2019] for $n_{1}$ rounds. Let $\widehat{\mathbf{\Theta}}_{m,n_{1}}$ be the estimate of $\mathbf{\Theta}_{m,*}$ at the end of $n_{1}$ rounds. Let the SVD of $\widehat{\mathbf{\Theta}}_{m,n_{1}}$ be given by $\operatorname{SVD}(\widehat{\mathbf{\Theta}}_{m,n_{1}})=\widehat{\mathbf{U}}_{m,n_{1}}\widehat{\mathbf{S}}_{m,n_{1}}\widehat{\mathbf{V}}_{m,n_{1}}$. Then LowOFUL rotates the actions as follows:
@@
Finally for $n_{2}=n-n_{1}$ rounds, LowOFUL invokes the specialized OFUL algorithm [Abbasi-Yadkori et al., 2011] for the rotated action set $\widetilde{\mathcal{A}}_{m}$ with the low dimension $k=\left(d_{1}+d_{2}\right)\kappa-\kappa^{2}$. Note that the LowOFUL runs the per-task low dimensional OFUL algorithm rather than learning the underlying structure across the tasks [Mukherjee et al., 2023].
@@
(1) The PreDeToR (-$\tau$) learns to exploit the underlying latent bilinear structure and reward correlation between the different tasks even when $A\geq n$ but $M_{\text{pre}}\gg A$.
(2) The PreDeToR (-$\tau$) has lower regret than its weak demonstrator Thomp which does not know the underlying bilinear structure.
(3) The PreDeToR (-$\tau$) has lower regret compared to DPT-greedy and AD. Also, observe that PreDeToR matches the performance of LowOFUL.
@@
Experimental Result: We observe these outcomes in Figure 3. In Figure 3(a) we experiment with rank $1$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure 3(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR learns the correlation of rewards across tasks and can perform well. Observe from Figure 3(a), and 3(b) that PreDeToR has lower regret than Thomp and matches LowOFUL. Also, in this low-data regime it is not enough for LowOFUL to learn the underlying $\mathbf{\Theta}_{m,*}$ with high precision. Hence, PreDeToR also has slightly lower regret than LowOFUL. The training of AD ensures that it does not outperform the demonstrator Thomp. Most importantly it shows that PreDeToR can exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
A.4 Empirical Study: Latent Bandits
In this section, we discuss the performance of PreDeToR (-$\tau$) against the other baselines in the latent bandit setting and create a generalized bilinear bandit setting. Note that the number of tasks $M_{\text{pre}}\gg A\geq n$. Using this experiment, we want to evaluate the ability of PreDeToR (-$\tau$) to exploit the underlying reward correlation when the horizon is small, the number of tasks is large, and understand its performance in the latent bandit setting [Hong et al., 2020, Maillard and Mannor, 2014, Pal et al., 2023, Kveton et al., 2017]. We create a latent bandit setting which generalizes the bilinear bandit setting [Jun et al., 2019, Lu et al., 2021, Kang et al., 2022, Mukherjee et al., 2023]. Again note that this setting also goes beyond the linear feedback model [Abbasi-Yadkori et al., 2011, Lattimore and Szepesvári, 2020] and is related to matrix bandits [Yang and Wang, 2020].
@@ A.4 Empirical Study: Latent Bandits
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LowOFUL. However, we now implement a special LowOFUL (stated in Section A.3) which has knowledge of the shared latent parameters $\mathbf{U}$, and $\mathbf{V}$. We call this the LowOFUL (oracle) algorithm. Therefore LowOFUL (oracle) has knowledge of the problem parameters in the latent bandit setting and hence the name.
@@
(1) The PreDeToR (-$\tau$) learns to exploit the underlying latent bilinear structure and reward correlation between the different tasks even when $A\geq n$ but $M_{\text{pre}}\gg A$.
(2) The PreDeToR (-$\tau$) has lower regret than its weak demonstrator Thomp which does not know the underlying latent structure. The PreDeToR (-$\tau$) has lower regret compared to DPT-greedy.
(3) The PreDeToR (-$\tau$) has regret that is closer to LowOFUL (oracle).
@@
Experimental Result: We observe these outcomes in Figure 4. In Figure 4(a) we experiment with rank $1$ hidden parameter $\bm{\theta}_{m,*}\bm{\theta}_{m,*}^{\top}$ and latent parameters $\mathbf{u}\mathbf{v}^{\top}$ shared across the tasks and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure 4(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. In Figure 4(c) we experiment with rank $3$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=300000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR (-$\tau$) learns the correlation of rewards across tasks and is able to perform well. Observe from Figure 4(a), 4(b), and 4(c) that PreDeToR has lower regret than Thomp and has regret closer to LowOFUL (oracle)which has access to the problem-dependent parameters. Hence. LowOFUL (oracle) outperforms PreDeToR (-$\tau$) in this setting. The training of AD ensures that it does not outperform the demonstrator Thomp. This shows that PreDeToR is able to exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
A.5 Understanding PreDeToR
In this section, we try to understand the behavior of PreDeToR and its ability to exploit the reward correlation across tasks under a linear multivariate Gaussian model. In this model, the hidden task parameter, $\bm{\theta}_{*}$, is a random variable drawn from a multi-variate Gaussian distribution [Bishop, 2006] and the feedback follows a linear model. We study this setting since we can estimate the Linear Minimum Mean Square Estimator (LMMSE) in this setting [Carlin and Louis, 2008, Box and Tiao, 2011]. This yields a posterior prediction for the mean of each action over all tasks on average, by leveraging the linear structure when $\bm{\theta}_{*}$ is drawn from a multi-variate Gaussian distribution. So we can compare the performance of PreDeToR against such an LMMSE and evaluate whether it is exploiting the underlying linear structure and the reward correlation across tasks.

### v2

@@ A.2.1 Experimental Details
Non-Linear Bandit: We now consider the setting when $f(\mathbf{x},\bm{\theta}_{*})=1/(1+0.5\cdot\exp(2\cdot\exp(-\mathbf{x}^{\top}\bm{\theta}_{*})))$. Again, here $\mathbf{x}\in\mathbb{R}^{d}$ is the action feature, and $\bm{\theta}_{*}\in\mathbb{R}^{d}$ is the hidden parameter. Note that this is different than the generalized linear bandit setting (Filippi et al., 2010; Li et al., 2017). Again for every experiment, we first generate tasks from $\mathcal{T}_{\text{pre}}$. Then we sample a fixed set of actions from $\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}/d\right)$ in $\mathbb{R}^{d}$ and this constitutes the features. Then for each task $m\in[M]$ we sample $\bm{\theta}_{m,*}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}/d\right)$ to produce the means $\mu(m,a)=1/(1+0.5\cdot\exp(2\cdot\exp(-\mathbf{x}(m,a)^{\top}\bm{\theta}_{m,*})))$ for $a\in\mathcal{A}$ and $m\in[M]$. Again note that in this setting $\mathbf{x}(m,a)$ for each $a\in\mathcal{A}$ is fixed for all tasks $m$.
@@
A.2.2 Details of Baselines
(1) Thomp: This baseline is the stochastic $A$-action bandit Thompson Sampling algorithm from Thompson (1933); Agrawal & Goyal (2012); Russo et al. (2018); Zhu & Tan (2020). We briefly describe the algorithm below: At every round $t$ and each action $a$, Thomp  samples $\gamma_{m,t}(a)\sim\mathcal{N}(\widehat{\mu}_{m,t-1}(a),\sigma^{2}/N_{m,t-1}(a))$, where $N_{m,t-1}(a)$ is the number of times the action $a$ has been selected till $t-1$, and $\widehat{\mu}_{m,t-1}(a)=\frac{\sum_{s=1}^{t-1}\widehat{r}_{m,s}\mathbf{1}(I_{s}=a)}{N_{m,t-1}(a)}$ is the empirical mean. Then the action selected at round $t$ is $I_{t}=\mathop{\rm arg\,max}_{a}\gamma_{m,t}(a)$. Observe that Thomp is not a deterministic algorithm like UCB (Auer et al., 2002). So we choose Thomp as the weak demonstrator $\pi^{w}$ because it is more exploratory than UCB and also chooses the optimal action, $a_{m,*}$, a sufficiently large number of times. Thomp is a weak demonstrator as it does not have access to the feature set $\mathcal{X}$ for any task $m$.
(3) MLinGreedy: This is the multi-task linear regression bandit algorithm proposed by Yang et al. (2021). This algorithm assumes that there is a common low dimensional feature extractor $\mathbf{B}\in\mathbb{R}^{k\times d}$, $k\leq d$ shared between the tasks and the rewards per task $m$ are linearly dependent on a hidden parameter $\bm{\theta}_{m,*}$. Under a diversity assumption (which may not be satisfied in real data) and $\mathbf{W}=\left[\mathbf{w}_{1},\ldots,\mathbf{w}_{M}\right]$ they assume $\mathbf{\Theta}=\left[\bm{\theta}_{1,*},\ldots,\bm{\theta}_{M,*}\right]=\mathbf{B}\mathbf{W}$. During evaluation MLinGreedy estimates the $\mathbf{\widehat{B}}$ and $\widehat{\mathbf{W}}$ from training data and fit $\widehat{\bm{\theta}}_{m}=\mathbf{\widehat{B}}\widehat{\mathbf{w}}_{m}$ per task and selects action greedily based on $I_{m,t}=\mathop{\rm arg\,max}_{a}\mathbf{x}_{m,a}^{\top}\widehat{\bm{\theta}}_{m,*}$. Finally, note that MLinGreedy requires access to the action features to estimate $\widehat{\bm{\theta}}_{m}$ and select actions as opposed to DPT, AD, and PreDeToR.
A.3 Empirical Study: Comparison against K-armed bandits and DPT
In this section, we discuss the performance of PreDeToR (-$\tau$) when there is no latent structure in the data, that is the $K$-armed bandits. Then we compare the performance of PreDeToR (-$\tau$) against DPT.
Baselines: In the K-armed bandits We implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB. In the linear and non-linear setting, we compare against DPT instead of DPT-greedy.
Settings: In the K-armed bandit setting we consider $d=6$, and the arms as canonical vectors $\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{6}$. For each task $m$, we choose the hidden parameter $\bm{\theta}_{m,*}$ similar to the linear bandit setting discussed in Section 5. Note that this results in a K-armed bandit setting. For the linear and non-linear setting comparison, we use the same setting as Section 5, and 4.
Outcomes: We first discuss the main outcomes from our experimental results in K-armed bandits and then in comparison against DPT in linear and non-linear settings.
Finding 7: PreDeToR (-$\tau$) matches the performance of the demonstrator when there is no structure (K-armed bandits). PreDeToR (-$\tau$) performs close to DPT in the linear and non-linear setting showing the usefulness of learning the reward structure.
(a) K-armed Bandit
(b) Comparison against DPT in linear setting
(c) Comparison against DPT in non-linear setting
Figure 7: Experiment with k-armed bandits and DPT (original). The y-axis shows the cumulative regret.
Experimental Result: We observe these outcomes in Figure 7. In Figure 7(a) the demonstrator $\pi^{w}$ is the Thomp algorithm. Note that there is no structure across arms now, and sampling one arm gives no information about other arms in a task. We observe that PreDeToR-$\tau$ performs similarly to the demonstrator Thomp, and also shows that incorporating exploration is a sound technique. Also, AD performs similarly to the demonstrator Thomp. Both DPT-greedy and PreDeToR fail to learn the latent structure across the tasks and therefore do not learn any exploration strategy.
In Figure 7(b) we show the linear bandit setting discussed in Section A.2. We observe that PreDeToR (-$\tau$) matches the performance of DPT, and LinUCB. Note that DPT has access to the optimal action per task, and LinUCB is the optimal oracle algorithm that leverages the structure information.
In Figure 7(c) we show the non-linear bandit setting discussed in Section A.2. Again we observe that PreDeToR (-$\tau$) matches the performance of DPT and has lower cumulative regret than AD and LinUCB which fails to perform well in this non-linear setting due to its algorithmic design.
A.4 Empirical Study: Bilinear Bandits
@@
In this section, we discuss the performance of PreDeToR against the other baselines in the bilinear setting. Again note that the number of tasks $M_{\text{pre}}\gg A\geq n$. Through this experiment, we want to evaluate the performance of PreDeToR to exploit the underlying latent structure and reward correlation when the horizon is small, the number of tasks is large, and understand its performance in the bilinear bandit setting (Jun et al., 2019; Lu et al., 2021; Kang et al., 2022; Mukherjee et al., 2023). Note that this setting also goes beyond the linear feedback model (Abbasi-Yadkori et al., 2011; Lattimore & Szepesvári, 2020) and is related to matrix bandits (Yang & Wang, 2020).
@@
This special case is studied in Chaudhuri et al. (2017).
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, and Thomp. Note that we do not implement the LinUCB and MLinGreedy for the bilinear bandit setting. However, we now implement the LowOFUL (Jun et al., 2019) which is optimal in the bilinear bandit setting.
LowOFUL: The LowOFUL algorithm first estimates the unknown parameter $\mathbf{\Theta}_{m,*}$ for each task $m$ using E-optimal design (Pukelsheim, 2006; Fedorov, 2013; Jun et al., 2019) for $n_{1}$ rounds. Let $\widehat{\mathbf{\Theta}}_{m,n_{1}}$ be the estimate of $\mathbf{\Theta}_{m,*}$ at the end of $n_{1}$ rounds. Let the SVD of $\widehat{\mathbf{\Theta}}_{m,n_{1}}$ be given by $\operatorname{SVD}(\widehat{\mathbf{\Theta}}_{m,n_{1}})=\widehat{\mathbf{U}}_{m,n_{1}}\widehat{\mathbf{S}}_{m,n_{1}}\widehat{\mathbf{V}}_{m,n_{1}}$. Then LowOFUL rotates the actions as follows:
@@
Finally for $n_{2}=n-n_{1}$ rounds, LowOFUL invokes the specialized OFUL algorithm (Abbasi-Yadkori et al., 2011) for the rotated action set $\widetilde{\mathcal{A}}_{m}$ with the low dimension $k=\left(d_{1}+d_{2}\right)\kappa-\kappa^{2}$. Note that the LowOFUL runs the per-task low dimensional OFUL algorithm rather than learning the underlying structure across the tasks (Mukherjee et al., 2023).
@@
Finding 8: PreDeToR (-$\tau$) outperforms DPT-greedy, AD, and matches the performance of LowOFUL in bilinear bandit setting.
@@
Experimental Result: We observe these outcomes in Figure 8. In Figure 8(a) we experiment with rank $1$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure 8(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR learns the correlation of rewards across tasks and can perform well. Observe from Figure 8(a), and 8(b) that PreDeToR has lower regret than Thomp and matches LowOFUL. Also, in this low-data regime it is not enough for LowOFUL to learn the underlying $\mathbf{\Theta}_{m,*}$ with high precision. Hence, PreDeToR also has slightly lower regret than LowOFUL. Note that the main objective of AD is to match the performance of its demonstrator. Most importantly it shows that PreDeToR can exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
A.5 Empirical Study: Latent Bandits
In this section, we discuss the performance of PreDeToR (-$\tau$) against the other baselines in the latent bandit setting and create a generalized bilinear bandit setting. Note that the number of tasks $M_{\text{pre}}\gg A\geq n$. Using this experiment, we want to evaluate the ability of PreDeToR (-$\tau$) to exploit the underlying reward correlation when the horizon is small, the number of tasks is large, and understand its performance in the latent bandit setting (Hong et al., 2020; Maillard & Mannor, 2014; Pal et al., 2023; Kveton et al., 2017). We create a latent bandit setting which generalizes the bilinear bandit setting (Jun et al., 2019; Lu et al., 2021; Kang et al., 2022; Mukherjee et al., 2023). Again note that this setting also goes beyond the linear feedback model (Abbasi-Yadkori et al., 2011; Lattimore & Szepesvári, 2020) and is related to matrix bandits (Yang & Wang, 2020).
@@ A.5 Empirical Study: Latent Bandits
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LowOFUL. However, we now implement a special LowOFUL (stated in Section A.4) which has knowledge of the shared latent parameters $\mathbf{U}$, and $\mathbf{V}$. We call this the LowOFUL (oracle) algorithm. Therefore LowOFUL (oracle) has knowledge of the problem parameters in the latent bandit setting and hence the name. Again note that we do not implement the LinUCB and MLinGreedy for the latent bandit setting.
@@
Finding 9: PreDeToR (-$\tau$) outperforms DPT-greedy, AD, and matches the performance of LowOFUL (oracle) in latent bandit setting.
@@
Experimental Result: We observe these outcomes in Figure 9. In Figure 9(a) we experiment with rank $1$ hidden parameter $\bm{\theta}_{m,*}\bm{\theta}_{m,*}^{\top}$ and latent parameters $\mathbf{u}\mathbf{v}^{\top}$ shared across the tasks and set horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=30$, and $d=5$. In Figure 9(b) we experiment with rank $2$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=250000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. In Figure 9(c) we experiment with rank $3$ hidden parameter $\mathbf{\Theta}_{m,*}$, and latent parameters $\mathbf{U}\mathbf{V}^{\top}$ and set horizon $n=20$, $M_{\text{pre}}=300000$, $M_{\text{test}}=200$, $A=25$, and $d=5$. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, AD and Thomp. Note that for any task $m$ for the horizon $20$ the Thomp will be able to sample all the actions at most once. Note that for this small horizon setting the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. In contrast PreDeToR (-$\tau$) learns the correlation of rewards across tasks and is able to perform well. Observe from Figure 9(a), 9(b), and 9(c) that PreDeToR has lower regret than Thomp and has regret closer to LowOFUL (oracle)which has access to the problem-dependent parameters. Hence. LowOFUL (oracle) outperforms PreDeToR (-$\tau$) in this setting. This shows that PreDeToR is able to exploit the underlying latent structure and reward correlation better than DPT-greedy, and AD.
A.6 Connection between PreDeToR and Linear Multivariate Gaussian Model
In this section, we try to understand the behavior of PreDeToR and its ability to exploit the reward correlation across tasks under a linear multivariate Gaussian model. In this model, the hidden task parameter, $\bm{\theta}_{*}$, is a random variable drawn from a multi-variate Gaussian distribution (Bishop, 2006) and the feedback follows a linear model. We study this setting since we can estimate the Linear Minimum Mean Square Estimator (LMMSE) in this setting (Carlin & Louis, 2008; Box & Tiao, 2011). This yields a posterior prediction for the mean of each action over all tasks on average, by leveraging the linear structure when $\bm{\theta}_{*}$ is drawn from a multi-variate Gaussian distribution. So we can compare the performance of PreDeToR against such an LMMSE and evaluate whether it is exploiting the underlying linear structure and the reward correlation across tasks. We summarize this as follows:
Finding 10: PreDeToR learns the reward correlation covariance matrix from the in-context training data $\mathcal{H}_{\text{train}}$ and acts greedily on it.